A rocket speeding by the Earth is known, when at rest, to be roughly cylindrical with a radius of 14 m and a length of 140 m. When speeding by, its volume is measured to be 3.83 × 104 m3. What is its speed?

Insufficient information is given.

0.896c

0.774c.

0.678c
Answer #1

When the rocket is speeding, the change in length occurs for that length only which is in direction of motion. Since the length of cylinder is in the direction of motion hence only the there will be change in the length of cylinder and not in the radius of cylinder.

now using the relation

here L is final length and l is initial length, v is velocity of rocket and c is speed of light

Since Area x length = volume of cylinder

hence

hence 3.83 x 10^4 = 3.14 * 14^2 * 140 *

c = 3 x 10^8 m/s

v/c = 0.77

hence v = 0.77c therefore option 2 is correct